### v1

@@ 4.2 Subject Causal Fingerprinting
We first verify the effectiveness of the proposed Causal-Dynamics Model combined with Modal Decomposition and Projection (CM+MD&P) for subject fingerprinting. Previous research has established that resting-state fMRI shows considerable differences in brain activation patterns among individuals [57]. The dataset consists of two resting sessions, Rest1 and Rest2, each with two scanning orders, LR and RL. We select one of the four possible combinations to create a labeled database, denoted as $\mathcal{R}_{h}$, and use the remaining three combinations as query data.
[…]
@@
Method CM+GNN CM+RF CM+SVM FC+GNN R+GNN R+DNN R+DBRNN R+BAnD
Rest 100% 100% 99.232% 97.307% 96.629% 95.164% 94.355% 100%
Emot 99.872% 99.488% 85.677% 89.160% 69.238% 72.912% 84.757% 98.493%
Gamb 96.122% 74.168% 93.094% 75.553% 81.683% 68.535% 69.023% 94.176%
Lang 99.159% 96.675% 89.514% 85.755% 89.177% 81.517% 82.204% 97.491%
Moto 96.733% 80.562% 96.675% 77.982% 76.858% 66.374% 77.773% 95.754%
Rela 98.051% 81.419% 81.841% 90.609% 86.554% 87.700% 79.628% 93.121%
Soci 97.544% 75.703% 77.749% 69.811% 73.706% 71.455% 69.950% 96.385%
WMem 99.523% 90.767% 95.396% 78.760% 80.202% 76.620% 82.446% 98.299%
Average 98.376% 87.350% 89.897% 83.117% 81.756% 78.929% 80.017% 96.710%
@@
Additionally, we note that the BAnD method [44] leverages volumetric data and varying cohort sizes within an end-to-end deep learning framework. Although volumetric data offers higher resolution than the Schaefer-100 parcellation used in our study, Raw+BAnD attains a slightly lower average accuracy (96.710%) compared to CM+GNN. This result emphasizes the benefit of integrating causal signatures with GNN-based architectures.
To further assess the reliability of our causal-based approach, we tested it on an independent dataset [60]. The method achieved nearly 95% accuracy in task fingerprinting, confirming its consistency and wide applicability.
@@ 5.1 Directed Causal Interaction
Figure 8: Reachability-based heatmaps for one subject across eight cognitive tasks (Rest, Emot, Gamb, Lang, Moto, Rela, Soci, and WMem). Brighter colors (yellow) indicate stronger maximum neural activations, while darker tones (blue) reflect lower excitability. Red dashed circles highlight representative activation hotspots corresponding to critical brain regions known for each cognitive task. These patterns highlight functional specialization differences across diverse cognitive tasks.
@@ 6 Conclusion
Despite the novelty of the proposed approach, we remark on the following limitations: i) In our approach, system inputs were manually specifically selected based on different functional brain regions, while other ways of choosing system state input regions may potentially further improve the performance. ii) While our subject fingerprinting method is model-based and generalizable without training, the task fingerprinting method relies on training a Graph Neural Network (GNN), which may limit its generalization across datasets.
@@
In future work, we plan to: i) Explore causal fingerprints and the reachability landscape concept in both healthy controls (e.g., through other open neuroimaging datasets) and in cases of neurodegenerative diseases such as Alzheimer’s disease [52, 11, 70]; ii) Enhance the generalizability of our task fingerprinting approach, potentially through transfer learning techniques to accelerate adaptation to new datasets.
@@ 7 Acknowledgement
8 Disclaimer
“The views expressed in this manuscript are those of the author(s) and do not reflect the official policy or position of the U.S. Naval Academy, Department of the Navy, the Department of Defense, or the U.S. Government.”
@@ 8 Disclaimer
We train the GNN in a supervised fashion using all $R_{s,h}^{D}$ labeled by task $h$. Each node in the graph corresponds to one brain region, whose feature vector is that region’s row in $R_{s,h}$. The normalized $A$ block from $R_{s,h}$ sets up the adjacency. We minimize a negative log-likelihood loss comparing the GNN’s predicted task label to the true $h$. To enhance generalization and prevent overfitting, we employ weight decay (L2 regularization) within AdamW, in tandem with SAM [71] as a meta-optimizer. Specifically, In each training iteration, SAM and AdamW operate in concert. AdamW is the base optimizer, handling parameter updates with weight decay, while SAM adds an extra “perturbation” step to reduce the impact of sharp minima. During training, we reset the learning rate every $T_{0}=20$ epochs using a cosine decay schedule. Each cycle doubles in length via $T_{\mathrm{mult}}=2$, and the minimum learning rate is set to $\eta_{\min}=10^{-6}$. This combination leverages weight decay, sharpness‐aware optimization, and dynamic learning‐rate control to yield robust parameter solutions, even for noisy or short fMRI time series. At inference time, a query $R^{Q}$ is similarly fed into the GNN, producing a final log‐softmax distribution over tasks. The highest‐scoring task index is selected as our predicted label. Through this causal‐based GNN approach, we effectively leverage the multi‐timescale structure of $\bigl{[}Q\,A\,B_{1}\,B_{2}\bigr{]}$ to classify fMRI tasks with high accuracy.

### v2

@@ 4.2 Subject Causal Fingerprinting
We first verify the effectiveness of the proposed Causal-Dynamics Model combined with Modal Decomposition and Projection (CM+MD&P) for subject fingerprinting. Previous research has established that resting-state fMRI shows considerable differences in brain activation patterns among individuals [65]. The dataset consists of two resting sessions, Rest1 and Rest2, each with two scanning orders, LR and RL. Preprocessing follows Section 4.1.
We run four cross-condition folds: each of the four resting scans (Rest1-LR, Rest1-RL, Rest2-LR, Rest2-RL) is taken in turn as the reference database, denoted as $\mathcal{R}_{h}$, while the remaining three scans serve as queries. This guarantees that reference and query data are always disjoint. Within each fold, model parameters $(Q,A,B_{1},B_{2})$ are estimated only from the reference run; the ensuing modal matrices serve as one-shot “causal signatures” and are not updated during querying. Both baselines and the proposed method use the same four‐fold split and identical preprocessing.
The subject fingerprint is obtained by solving (11). For the results presented below, we chose the feature set as $\mathcal{T}=[\widehat{T}$], which relies only on the dynamic models of slow interaction and ignores the fast one. More details about this will be discussed in Remark 2. The subject fingerprint result, presented in Table 1, demonstrate that the causal fingerprinting method, specifically CM+MD&P, consistently achieves an accuracy rate above 94% for all resting sessions. The highest accuracy observed is 96.334% for the Rest2 LR session. This performance significantly surpasses that of the correlation-based baseline method, FC+CoR, which averages approximately 45% accuracy. Even though capturing causal interactions is harder, our method not only outperforms the FC+CoR baseline but also other alternative approaches such as CM+CoR, CM+FN, and FC+MD&P. Tangent-Space Fingerprinting (TSF) achieves $88\%$ accuracy, about 66  percentage points behind CM+MD&P, because it models only the Riemannian geometry of static functional connectivity. The PH-H0 landscape method attains $85\%$; its zero-dimensional topological summary likewise ignores directionality and temporal scale, leaving it $10$  percentage points short of our causal approach. To ensure a fair comparison, we applied the CM+MD&P method as described in Sec. 3.2 and the FC+CoR method as outlined in the 2015 study by FES et al. [18] directly to our dataset, without any further adjustments or preprocessing. Although additional data processing could potentially improve the accuracy of all methods tested, our main goal is to confirm the viability and efficacy of causal signatures for individual identification. Therefore, we have deliberately avoided such optimizations to maintain an impartial comparison. To assess the generalizability of our approach, we ran experiments on a public dataset [66]. This dataset contains resting‑state fMRI scans from healthy controls and Alzheimer’s disease patients. It does not include repeated sessions that match our original design. To work around this, we split one long resting‑state scan into two separate segments. This produced two distinct datasets for analysis. Our method achieved nearly 99% accuracy in subject fingerprinting. These results highlight its strong performance and replicability across different populations and scanning protocols.
@@
Method CM+GNN CM+RF CM+SVM FC+GNN R+GNN R+DNN R+DBRNN R+BAnD R+TCN–BiLSTM
Rest 100% 100% 99.232% 97.307% 96.629% 95.164% 94.355% 100% 99.532%
Emot 99.872% 99.488% 85.677% 89.160% 69.238% 72.912% 84.757% 98.493% 93.445%
Gamb 96.122% 74.168% 93.094% 75.553% 81.683% 68.535% 69.023% 94.176% 89.395%
Lang 99.159% 96.675% 89.514% 85.755% 89.177% 81.517% 82.204% 97.491% 91.304%
Moto 96.733% 80.562% 96.675% 77.982% 76.858% 66.374% 77.773% 95.754% 88.992%
Rela 98.051% 81.419% 81.841% 90.609% 86.554% 87.700% 79.628% 93.121% 89.663%
Soci 97.544% 75.703% 77.749% 69.811% 73.706% 71.455% 69.950% 96.385% 92.118%
WMem 99.523% 90.767% 95.396% 78.760% 80.202% 76.620% 82.446% 98.299% 94.327%
Average 98.376% 87.350% 89.897% 83.117% 81.756% 78.929% 80.017% 96.710% 92.347%
@@
TCN–BiLSTM: Temporal Convolutional Network–Bidirectional Long Short–Term Memory [45]
@@
We also evaluated the R + TCN–BiLSTM baseline, a Temporal Convolutional Network followed by bidirectional LSTM layers trained end to end on raw fMRI series [45]. Its mean accuracy of $92.3\%$ is higher than the other raw‑data deep nets (R + GNN, R + DNN, R + DBRNN) yet remains about $6$ percentage points below CM + GNN. While the convolutional blocks capture local temporal motifs, the model lacks directed and multi‑timescale information, limiting its discriminative power relative to our causal signatures.
To further assess the reliability of our causality-based approach, we tested it on an independent dataset [68]. The method achieved nearly 95% accuracy in task fingerprinting, confirming its consistency and wide applicability.
4.4 Effect of sampling interval $\Delta t$ on model fingerprinting accuracy
Table 4: Subject fingerprinting accuracy (%) under different sampling intervals $\Delta t$ (seconds)
$\Delta t$ (s) $\mathbf{\Delta t=0.72}$ $\mathbf{\Delta t=1.44}$ $\mathbf{\Delta t=2.16}$
Rest1 LR 95.396% 90.143% 80.636%
Rest2 LR 96.334% 92.002% 81.452%
Rest1 RL 94.970% 88.747% 78.890%
Rest2 RL 94.203% 89.563% 77.938%
Note that the sampling interval $\Delta t$ is critical to our discrete-time causal model. In this experiment, we adopt the native sampling interval $\Delta t=0.72$s because it is the scanner’s repetition time (TR), as specified in the data description [61]. Repetition time is the time interval between the beginning of one MRI pulse sequence and the beginning of the next for the same slice. Staying at this original resolution avoids any resampling artifacts and preserves the rapid hemodynamic changes our two-timescale model is meant to capture.
Nevertheless, it is still important to discuss the sensitivity of our algorithm to the sampling interval $\Delta t$. Table 4 summarizes how resampling affects subject fingerprinting. Accuracy is computed for each of the four resting runs at three resolutions: the native $0.72$s, and down‑sampled versions at $1.44$s (every second frame) and $2.16$s (every third frame).
The results in Table 4 confirm that accuracy peaks at the native interval ($\Delta t=0.72$s), the scanner’s inherent repetition time in the HCP protocol and therefore the finest temporal resolution available. Accuracy drops by roughly five percentage points when the data are down-sampled to $1.44$ s and falls a further ten points at $2.16$ s. The two-timescale model evidently relies on sub-second variations that are smoothed out at coarser resolutions, degrading the estimated matrices Q&AQ\&A blocks. We therefore retain $\Delta t=0.72$ s for all reported experiments to preserve temporal fidelity while avoiding unnecessary information loss.
Task fingerprinting is less sensitive: accuracy falls from $98.376\%$ at $\Delta t=0.72$ s to $96.442\%$ at $1.44$ s and $92.127\%$ at $2.16$ s, suggesting that large-scale task-evoked activation patterns remain highly discriminative even when the temporal resolution is moderately reduced.
@@ 5.1 Directed Causal Interaction
Figure 8: Reachability-based heatmaps for one subject across eight cognitive tasks (Rest, Emot, Gamb, Lang, Moto, Rela, Soci, and WMem). Brighter colors (yellow) indicate stronger maximum neural activations, while darker tones (blue) reflect lower excitability. Red dashed circles highlight representative activation hotspots corresponding to critical brain regions known for each cognitive task. The black boundary outlines the edge of the flattened cortical map. The correspondence between this reachability representation and anatomical brain regions is illustrated in Fig. 4. These patterns highlight functional specialization differences across diverse cognitive tasks.
@@ 6 Conclusion
Despite the novelty of the proposed approach, we remark on the following limitations: i) In our approach, system inputs were manually specifically selected based on different functional brain regions, while other ways of choosing system state input regions may potentially further improve the performance. ii) While our subject fingerprinting method is model-based and generalizable without training, the task fingerprinting method relies on training a Graph Neural Network (GNN), which may limit its generalization across datasets. iii) Subject fingerprinting shows measurable sensitivity to the sampling interval $\Delta t$ (coarser sampling degrades accuracy).
@@
In future work, we plan to: i) Explore causal fingerprints and the reachability landscape concept in both healthy controls (e.g., through other open neuroimaging datasets) and in cases of neurodegenerative diseases such as Alzheimer’s disease [55, 11, 78]; ii) Enhance the generalizability of our task fingerprinting approach, potentially through transfer learning techniques to accelerate adaptation to new datasets. iii) Systematically assess robustness to choices of sampling interval $\Delta t$, parcellation, and preprocessing, and report performance in a manner that is independent of the acquisition and processing pipeline; iv) Extend the two-timescale linear model to piecewise linear or mildly nonlinear dynamics with uncertainty quantification, and replace manual selection of input regions with a data-driven procedure to reduce design bias.
@@ 7 Acknowledgement
[…]
Subject – CM+MD&P vs. baselines
[…]
Task – CM+GNN vs. baselines
[…]
We train the GNN in a supervised fashion using all $R_{s,h}^{D}$ labeled by task $h$. Each node in the graph corresponds to one brain region, whose feature vector is that region’s row in $R_{s,h}$; the normalised $A$ block sets up the adjacency. We minimise a negative log-likelihood loss comparing the GNN’s predicted task label to the true $h$. To enhance generalisation and prevent over-fitting, we employ weight decay (L2 regularisation) within AdamW, in tandem with SAM [79] as a meta-optimiser. Specifically, in each training iteration SAM and AdamW operate in concert: AdamW (learning rate $1\times 10^{-3}$, weight decay $5\times 10^{-4}$) performs the parameter update, while SAM adds a perturbation step of radius $\rho=0.02$ to avoid sharp minima. The learning rate is reset every $T_{0}=20$ epochs by a cosine warm-restart schedule; each cycle length doubles via $T_{\mathrm{mult}}=2$ and decays to $\eta_{\min}=10^{-6}$. We use mixed precision (AMP), gradient clipping at $1.0$, mini-batches of size $32$, and early stopping with patience $20$. On-the-fly data augmentation further improves robustness: Gaussian feature noise with $\sigma=0.1$, random edge drop of $20\%$, and subgraph sampling that keeps $80\%$ of the nodes. At inference time, a query $R^{Q}$ is fed into the same pipeline, producing a log-softmax distribution over tasks; the highest score is taken as the predicted label. Through this causal-based GNN we leverage the multi-timescale structure of $\bigl[Q\,A\,B_{1}\,B_{2}\bigr]$ to classify fMRI tasks with high accuracy.
Results Validation
For each baseline we construct a length-$391$ vector of subject‑wise accuracies (one paired value per subject, averaged over the four cross‑condition folds) and apply a two‑sided paired tt‑test after verifying normality with the Shapiro–Wilk test. Under a Bonferroni adjustment for the seven subject‑level comparisons ($\alpha=0.05/7\approx 0.0071$), every contrast with the proposed CM+MD&P remains highly significant, with pp‑values spanning $4.7\times 10^{-22}$ to $10^{-33}$ (Table 5, left).
For task fingerprinting we test eight baselines, so the corrected threshold is $\alpha=0.05/8\approx 0.0063$. All baselines, including the newly added R + TCN–BiLSTM, are significantly different from the proposed CM+GNN except R + BAnD, whose $p=1.6\times 10^{-2}$ exceeds the threshold (Table 5, right).
Mean accuracies with $95\,\%$ confidence intervals across folds are $95.4\%\pm 0.3$ for subject fingerprinting and $98.4\%\pm 0.2$ for task fingerprinting, indicating that the estimates are stable. Ten random perturbations of the reference/query split alter subject accuracy by no more than $0.6$ percentage points (standard deviation), confirming that the results are not driven by a particular split.